OpenStudy (anonymous):

Find the product. (4x-y)^3

OpenStudy (anonymous):

Use the formula \[ (a-b)^3=a^3-3 a^2 b+3 a b^2-b^3 \]

OpenStudy (anonymous):

a=4x b=y

OpenStudy (anonymous):

You will get \[ (4 x-y)^3=64 x^3-48 x^2 y+12 x y^2-y^3 \]